OpenStudy (anonymous):

convert 3x-4y=4 into slope-intercept form

OpenStudy (anonymous):

\[ 3x-4y=4 \implies 3x-4 = 4y \implies y = \frac {3x - 4} {4} \implies y = \frac34x - 1 \]